# Mitigate the Gap: Investigating Approaches for Improving Cross-Modal Alignment in CLIP

### v2

@@ Appendix B Limitations
For the zero-shot classification experiments, we have tested the models using the ImageNet-1k, CIFAR-10, CIFAR-100, Stanford Cars, and Food-101 datasets. More classification analysis using other benchmarks datasets such as Flower-102 and Hatefule Memes has not been studied in this work. Furthermore, this study is limited to the English language and further analyses on multilingual and cross-lingual representation learning are necessary. Moreover, the re-scaling mechanism in the intra-modality separation loss is dependant on the choice of pre-trained sentence encoder and has not been fully benchmarked in this work.

### v3

@@ Appendix B Limitations
For the zero-shot classification experiments, we have tested the models using the ImageNet-1k, CIFAR-10, CIFAR-100, Stanford Cars, and Flowers-102 datasets. More classification analysis using other benchmarks datasets such as Food-100 and Hateful Memes has not been studied in this work. Furthermore, this study is limited to the English language and further analyses on multilingual and cross-lingual representation learning are necessary. Moreover, the re-scaling mechanism in the intra-modality separation loss is dependant on the choice of the pre-trained sentence encoder and has not been fully benchmarked in this work.